Given the following information, calculate the expected return for the portfolio of stocks A and B. (Answer to the nearest tenth of a percent, but do not use a percent sign). Amount Expected Invested Return Stock A $3,000 19% Stock B $6,000 15%
Answer #1

Total investment=(3000+6000)=9000

Portfolio return=Respective return*Respective investment weight

=(3000/9000*19)+(6000/9000*15)

=16.3(Approx).